# Self-consistency error correction for accurate machine learning potentials from variational Monte Carlo

### v1

@@ 3.1 Correcting the SCE bias in the Hugoniot dataset
To assess the magnitude of the SCE in the Hugoniot dataset, we use a JSD WF with a basis set of [4s2p1d] and [2s2p1d] GTOs for the antisymmetric part and Jastrow factor, respectively. The same primitive basis set was used in Ref. 20. We initialize the SD by running an LDA-DFT calculation with twisted boundary conditions at the Baldereschi point, i.e, at $\mathbf{k}=\left(\frac{1}{4},\frac{1}{4},\frac{1}{4}\right)$ in crystal units. The Jastrow part is optimized using the stochastic reconfiguration method 50.
@@
We observe a significant bias for both quantities, with a discrepancy as large as $\sim 10\%$ of their values. This clearly highlights the importance of curing the SCE for the biased dataset. For this reason, we applied the corrections of Eqs. (10) and (11) to forces and pressures, respectively. For forces, we computed the numerical derivatives of the parameters $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}}$ with the FDM from DFT-LDA, the theory used to obtain $\lambda_{k,k^{\prime}}$, with a displacement $\Delta R=\pm 0.003$. The derivatives of $\lambda_{k,k^{\prime}}$ with respect to the volume, entering in the pressure correction, were obtained with the FDM using relative volume variations of ±$\pm 0.03\%$. In Fig. 1 we observe that the corrected forces and pressure are perfectly compatible with their values directly computed through the fit of the VMC PES. Fig. 2 shows the values of both the biased and corrected force components (pressures) versus the forces (pressures) estimated from the PES, for several configurations. Notice how the pressure correction acts almost like a rigid shift of $\sim 1$ GPa. The root mean squared error (RMSE) of biased and corrected quantities is also shown, further demonstrating the effectiveness of the correction in removing the SCE.
@@
Sometimes, the added correction significantly increases the error bar $\sigma_{F}$ of the corrected forces. We then disregard the corrections that lead to forces with $0.0153\sigma_{F}>$ 0.015 Ha/Bohr. For pressures, the threshold value of $1.5\times 10^{-5}$ is applied to $3\sigma_{p}$. See App. A.2 for further details. In principle, the FDM used to estimate the parameter derivatives $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}}$ can be replaced with more robust and faster approaches (e.g., using linear response theory 51). We leave this possibility for future work.
@@
Figure 1: (a) Comparison among the biased pressure evaluated with an expression equivalent to Eq. (5) (red diamonds), the corrected pressure obtained by applying Eq. (11) (violet square), and the pressure obtained by fitting the PES (dashed green line and green diamond) for a selected configuration in the Hugoniot dataset20. The PES of the system is also shown (green dots and dash-dotted line). (b) Comparison among the biased force evaluated with Eq. (5) (red diamonds), the corrected force obtained by applying Eq. (10) (violet square), and the force calculated by fitting the PES (dashed green line and green diamond) for the $z$-component of the second hydrogen atom belonging to the same configuration as in (a). The PES of the system along the atomic displacement is also shown (green dots and dash-dotted line).
@@
Figure 2: (a) Value of the biased (red markers) and corrected (violet markers) pressure for $6$ different $128$-atom configurations as a function of the numerical pressure estimated from the derivative of the PES with respect to the volume. (b) Values of the biased (red markers) and corrected (violet markers) force components as a function of the numerical force estimated by the PES for one of the configurations belonging to the Hugoniot dataset20. The dashed line is a reference for perfect consistency.
@@ 3.2 Comparison between VMC datasets
To illustrate how using an SCE-corrected JSD dataset affects the MLIPs generation, we compare the performances of MLIPs trained on three different datasets: (i) the aforementioned "biased dataset" obtained with a JSD WF and with forces and pressure computed with the standard expressions (Eq. 5), (ii) a "corrected dataset" where the corrections of Eqs. (10) and (11) are applied as described in Sec. 2.2 and (iii) a "reference dataset", i.e., the one introduced in Ref. 20. In particular, the latter was obtained from an improved Jastrow-correlated AGP WF, where the Jastrow and antisymmetric part were both optimized, thanks to a combination of geminal embedded orbitals52 to reduce the number of variational parameters and a restricted optimization based on the locality of the AGP $\lambda_{k,k^{\prime}}$ matrix. The $\lambda_{k,k^{\prime}}$ coefficients were optimized for those basis set elements $k$ and $k^{\prime}$ centered on atoms closer than a cutoff distance of 4 Bohr radii20. No significant lowering of the variational energy was found beyond this threshold, suggesting that the dependence of the energy on the remaining parameters can be disregarded. Thanks to this approach, the SCE on this dataset was effectively removed and we thus decided to take the models trained on it as the reference. Notice how the explicit optimization of the whole WF, as the one performed in Ref. 20, is a procedure that is about an order of magnitude more expensive than the partial optimization of the frozen-orbital JSD WF, and requires advanced minimization techniques 53, 54.
For all the MLIPs, a KRR model is trained on the difference between VMC quantities and DFT ones, the latter using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional 55, following the $\Delta$-learning approach 18, 19, 20. In this way, a much higher accuracy is achieved with relatively small datasets such as the ones used here. The values of the hyperparameters chosen for these models are reported in App. A.1. Among the $558$ configurations of the datasets, we select $58$ configurations for testing. We investigate the effect of varying the training weights in the loss function (Eq. (3)) on the RMSE computed for energy, forces, and pressure of the test set for each dataset.
@@
$\Delta_{X}=\frac{\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{pred}})-\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{dummy}})}{\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{dummy}})},$ (12)
where $X_{\textrm{pred}}$ are the model predictions, $X_{\textrm{test}}$ the test set values, $E_{\textrm{dummy}}=\frac{1}{N_{test}}\sum_{i\in test}E_{i}$, $\mathbf{F}_{\textrm{dummy}}=0$, and $P_{\textrm{dummy}}=0$. In other words, $\Delta_{X}$ measures the relative improvement of the model prediction on the quantity $X$ with respect to a "dummy" model, corresponding to a perfectly flat PES equal to the average energy on the test set for all configurations. Notice how $\Delta_{X}=-1$ indicates that the model has learned exactly the quantity $X$, while a value $\Delta_{X}\geq 0$ means that the model has an error equal to, or larger than, a flat model. In the present work, the test quantities refer to the difference between the VMC and the DFT-PBE baseline, while $E_{\textrm{dummy}}$ is the average difference on the test set.
Figs. 3 and 4 show the results obtained for $\Delta_{E}$, $\Delta_{F}$, and $\Delta_{P}$ using MLIPs trained with different values of the training weights $W_{E}$, $W_{F}$, and $W_{P}$. We first analyze what happens by progressively increasing the weight on the forces for the different models without explicitly training on pressures (i.e., $W_{P}=0$). The force bias in the biased dataset is evident from the larger $\Delta_{F}$ of the corresponding model compared to that of both the corrected and reference datasets (see Fig. 3). Notice how, for small values of $W_{F}/W_{E}$, $\Delta_{F}$ is positive in the biased case, which in the $\Delta$-learning framework means that the correction on forces is detrimental. As the weight on forces is increased, all models reach a plateau in $\Delta_{F}$, while simultaneously losing some accuracy on the other quantities. For the energy, this loss of accuracy is significantly larger in the models trained on the biased dataset, further demonstrating the effect of the SCE.
@@
Finally, we mention how these conclusions exclusively depend on the quality of the datasets themself, i.e., their internal consistency, and not on the MLIPs specific architecture. To demonstrate this feature, in App. A.3, we discuss the relative improvement $\Delta_{X}$ for models obtained with the MACE 35 framework, which show the same behavior as the one found for the KRR MLIPs.
@@ 3.3 MD results comparison
To conclude our analysis, we run a set of MD simulations using MLIPs trained on the three different datasets. For each model, we select the relative training weights $W_{F}/W_{E}$ and $W_{P}/W_{E}$ to have a good tradeoff between the RMSE values on energy, forces, and pressure. In particular, for the MLIPs trained on the reference and corrected weight, we choose the last value of $W_{F}/W_{E}$ before the significant increase in $\Delta_{E}$ (i.e., the fourth smallest value in Fig. 3), in order to increase as much as possible the accuracy on forces without spoiling too much the energy. A similar reasoning is used to determine the optimal value of $W_{P}/W_{E}$ (see Fig. 4). For the MLIP trained on the biased dataset we choose a larger value of $W_{F}/W_{E}$, which is necessary to reach a reasonable accuracy on forces. This comes at the price of sacrificing the accuracy on the energy, a direct consequence of the SCE. The weights used for each MLIP are summarized in Tab. 1.
@@
We run MD simulations at three different density and temperature conditions close to the position of the deuterium Hugoniot curve given by the VMC-MLIP of Ref. 20. In particular, we considered thermodynamic conditions which span both the molecular liquid and the atomic one. Because of the $\Delta$-learning approach adopted, the resulting efficiency of these calculations is the same as a DFT Born-Oppeheimer MD simulation. Indeed, the energy, ionic forces and pressure of the system are calculated at each step with DFT, and the corresponding corrections are then added using the MLIP. Notice how the DFT baseline can be, in principle, replaced by a faster potential (e.g., using an MLIP trained on DFT data 56), improving the overall efficiency. Further details regarding the computational setup of the MD simulations are given in App. A.4.
A comparison of the equilibrium pressure during the dynamics for three different temperatures is reported in Tab. 2. Notice that, in principle, the reference dataset and the corrected one can give different results. Indeed, the optimization of the antisymmetric part of the WF not only improves the consistency of forces and pressure, but can also modify the PES of the system. Nevertheless, if compared with the biased model, the MLIP trained on the corrected data gives results that are consistently much closer to the reference ones. This suggests that the main discrepancy between the biased and reference models comes from the SCE, and that the SCE correction improves considerably the physical description of the system. We observe a maximum discrepancy of $\sim 2$ GPa at $r_{s}=2.02$ and $T=6000$ K between the biased and corrected results, while the corrected and reference ones are in statistical accordance within a joint error bar of 0.2 GPa at the same thermodynamic conditions. Remarkably, this state is very close to the Hugoniot compressibility peak, i.e., the maximum in the density reached by the shocked state along the Hugoniot curve. This thermodynamic point is very sensitive to changes in the underlying equation of state, and a pressure shift as the one observed here between the corrected and biased models has a significant impact in the final position of the Hugoniot 57. Similar conclusions can be reached by looking at the radial distribution function $g(r)$ (Fig. 5). Analogously to the pressure analysis, the $g(r)$ of the corrected model is compatible with the one given by the reference one. Also for this quantity the maximum discrepancy between the corrected and biased models is observed close to the compressibility peak at $T=6000$ K, which corresponds to the onset of the the molecular-to-atomic transition along the Hugoniot curve. These results further highlight how SCE forces and pressure corrections significantly improve the physical outcome of the resulting MLIPs.
@@ 4 Conclusions
Second, we showed how an easy-to-evaluate bias correction, previously developed in Ref. 22, can be applied to remove the SCE from the training set, and improve the accuracy of the resulting MLIPs.
@@
Our results reveal the importance of appropriately curing forces and pressures when they are computed with partially optimized WF in QMC and subsequently used to train MLIPs. The MLIPs quality degradation due to the SCE plaguing these VMC datasets is an effect that has been overlooked in previous applications. This could be even more relevant for heavier elements, since the SCE usually grows with the number of valence electrons 21. The availability of an SCE-correction framework used in combination with other approaches, such as $\Delta$-learning, to further increase the precision of the final models, opens the way to the efficient generation of accurate MLIP based on QMC electronic structure calculations.
@@ A.1 Kernel regression models parameters
In this Appendix we provide further details on the hyperparameters and specific kernels used for the models described in Sec. 3.2. The KRR models use a kernel as the one described in Ref. 19. Each environment is described by a density function written as a sum of Gaussian functions (with spread $\epsilon$) multiplied by a localized function $f_{c}$ depending on a cutoff radius $r_{c}$, i.e.,
@@
For the local energy (Eq. (4)) we also added an additional pair-wise term as done in Ref. 20, using cubic spline functions 58 defined on $10$ equally spaced grid points between $0.3$ Bohr and $5.5$ Bohr.

### v2

@@ 3.1 Correcting the SCE bias in the Hugoniot dataset
To assess the magnitude of the SCE in the Hugoniot dataset, we use a JSD WF with a basis set of [4s2p1d] and [2s2p1d] GTOs for the antisymmetric part and Jastrow factor, respectively. The same primitive basis set was used in Ref. Tenti2024. We initialize the SD by running an LDA-DFT calculation with twisted boundary conditions at the Baldereschi point, i.e, at $\mathbf{k}=\left(\frac{1}{4},\frac{1}{4},\frac{1}{4}\right)$ in crystal units. The Jastrow part is optimized using the stochastic reconfiguration method Sorella1998.
@@
We observe a significant bias for both quantities, with a discrepancy as large as $\sim 10\%$ of their values. This clearly highlights the importance of curing the SCE for the biased dataset. For this reason, we applied the corrections of Eqs. (10) and (11) to forces and pressures, respectively. For forces, we computed the numerical derivatives of the parameters $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}}$ with the FDM from DFT-LDA, the theory used to obtain $\lambda_{k,k^{\prime}}$, with a displacement $\Delta R=\pm 0.003$ Å. The derivatives of $\lambda_{k,k^{\prime}}$ with respect to the volume, entering in the pressure correction, were obtained with the FDM using relative volume variations of $\pm 0.03\%$. In Fig. 1 we observe that the corrected forces and pressure are perfectly compatible with their values directly computed through the fit of the VMC PES. Fig. 2 shows the values of both the biased and corrected force components (pressures) versus the forces (pressures) estimated from the PES, for several configurations. Notice how the pressure correction acts almost like a rigid shift of $\sim 1$ GPa. The root mean squared error (RMSE) of biased and corrected quantities is also shown, further demonstrating the effectiveness of the correction in removing the SCE.
@@
Sometimes, the added correction significantly increases the error bar $\sigma_{F}$ of the corrected forces. We then disregard the corrections that lead to forces with $3\sigma_{F}>0.015$ Ha/Bohr. For pressures, the threshold value of $1.5\times 10^{-5}$ is applied to $3\sigma_{p}$. See App. A.2 for further details. In principle, the FDM used to estimate the parameter derivatives $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}}$ can be replaced with more robust and faster approaches (e.g., using linear response theory Toulouse2018). We leave this possibility for future work.
@@
Figure 1: (a) Comparison among the biased pressure evaluated with an expression equivalent to Eq. (5) (red diamonds), the corrected pressure obtained by applying Eq. (11) (violet square), and the pressure obtained by fitting the PES (dashed green line and green diamond) for a selected configuration in the Hugoniot datasetTenti2024. The PES of the system is also shown (green dots and dash-dotted line). (b) Comparison among the biased force evaluated with Eq. (5) (red diamonds), the corrected force obtained by applying Eq. (10) (violet square), and the force calculated by fitting the PES (dashed green line and green diamond) for the $z$-component of the second hydrogen atom belonging to the same configuration as in (a). The PES of the system along the atomic displacement is also shown (green dots and dash-dotted line).
@@
Figure 2: (a) Value of the biased (red markers) and corrected (violet markers) pressure for $6$ different $128$-atom configurations as a function of the numerical pressure estimated from the derivative of the PES with respect to the volume. (b) Values of the biased (red markers) and corrected (violet markers) force components as a function of the numerical force estimated by the PES for one of the configurations belonging to the Hugoniot datasetTenti2024. The dashed line is a reference for perfect consistency.
@@ 3.2 Comparison between VMC datasets
To illustrate how using an SCE-corrected JSD dataset affects the MLIPs generation, we compare the performances of MLIPs trained on three different datasets: (i) the aforementioned "biased dataset" obtained with a JSD WF and with forces and pressure computed with the standard expressions (Eq. 5), (ii) a "corrected dataset" where the corrections of Eqs. (10) and (11) are applied as described in Sec. 2.2 and (iii) a "reference dataset", i.e., the one introduced in Ref. Tenti2024. In particular, the latter was obtained from an improved Jastrow-correlated AGP WF, where the Jastrow and antisymmetric part were both optimized, thanks to a combination of geminal embedded orbitalsSorella2015 to reduce the number of variational parameters and a restricted optimization based on the locality of the AGP $\lambda_{k,k^{\prime}}$ matrix. The $\lambda_{k,k^{\prime}}$ coefficients were optimized for those basis set elements $k$ and $k^{\prime}$ centered on atoms closer than a cutoff distance of 4 Bohr radiiTenti2024. No significant lowering of the variational energy was found beyond this threshold, suggesting that the dependence of the energy on the remaining parameters can be disregarded. Thanks to this approach, the SCE on this dataset was effectively removed and we thus decided to take the models trained on it as the reference. Notice how the explicit optimization of the whole WF, as the one performed in Ref. Tenti2024, is a procedure that is about an order of magnitude more expensive than the partial optimization of the frozen-orbital JSD WF, and requires advanced minimization techniques Holzmann2003, Slootman2024.
For all the MLIPs, a KRR model is trained on the difference between VMC quantities and DFT ones, the latter using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional Perdew1996, following the $\Delta$-learning approach Ramakrishnan2015, Tirelli2022, Tenti2024. In this way, a much higher accuracy is achieved with relatively small datasets such as the ones used here. The values of the hyperparameters chosen for these models are reported in App. A.1. Among the $558$ configurations of the datasets, we select $58$ configurations for testing. We investigate the effect of varying the training weights in the loss function (Eq. (3)) on the RMSE computed for energy, forces, and pressure of the test set for each dataset.
@@
$\Delta_{X}=\frac{\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{dummy}})-\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{pred}})}{\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{dummy}})},$ (12)
where $X_{\textrm{pred}}$ are the model predictions, $X_{\textrm{test}}$ the test set values, $E_{\textrm{dummy}}=\frac{1}{N_{test}}\sum_{i\in test}E_{i}$, $\mathbf{F}_{\textrm{dummy}}=0$, and $P_{\textrm{dummy}}=0$. In other words, $\Delta_{X}$ measures the relative improvement of the model prediction on the quantity $X$ with respect to a "dummy" model, corresponding to a perfectly flat PES equal to the average energy on the test set for all configurations. Notice how $\Delta_{X}=1$ when $\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{pred}})=0$, indicating that the model has learned exactly the quantity $X$, while a value $\Delta_{X}\leq 0$ means that the model has an error equal to, or larger than, a flat model. In the present work, the test quantities refer to the difference between the VMC and the DFT-PBE baseline, while $E_{\textrm{dummy}}$ is the average difference on the test set.
Figs. 3 and 4 show the results obtained for $\Delta_{E}$, $\Delta_{F}$, and $\Delta_{P}$ using MLIPs trained with different values of the training weights $W_{E}$, $W_{F}$, and $W_{P}$. We first analyze what happens by progressively increasing the weight on the forces for the different models without explicitly training on pressures (i.e., $W_{P}=0$). The force bias in the biased dataset is evident from the lower $\Delta_{F}$ of the corresponding model compared to that of both the corrected and reference datasets (see Fig. 3). Notice how, for small values of $W_{F}/W_{E}$, $\Delta_{F}$ is negative in the biased case, which in the $\Delta$-learning framework means that the correction on forces is detrimental. As the weight on forces is increased, all models reach a plateau in $\Delta_{F}$, while simultaneously losing some accuracy on the other quantities. For the energy, this loss of accuracy is significantly larger in the models trained on the biased dataset, further demonstrating the effect of the SCE.
@@
Finally, we mention how these conclusions exclusively depend on the quality of the datasets themself, i.e., their internal consistency, and not on the MLIPs specific architecture. To demonstrate this feature, in App. A.3, we discuss the relative improvement $\Delta_{X}$ for models obtained with the MACE Batatia2022 framework, which show the same behavior as the one found for the KRR MLIPs.
@@ 3.3 MD results comparison
To conclude our analysis, we run a set of MD simulations using MLIPs trained on the three different datasets. For each model, we select the relative training weights $W_{F}/W_{E}$ and $W_{P}/W_{E}$ to have a good tradeoff between the RMSE values on energy, forces, and pressure. In particular, for the MLIPs trained on the reference and corrected weight, we choose the last value of $W_{F}/W_{E}$ before the significant decrease in $\Delta_{E}$ (i.e., the fourth smallest value in Fig. 3), in order to increase as much as possible the accuracy on forces without spoiling too much the energy. A similar reasoning is used to determine the optimal value of $W_{P}/W_{E}$ (see Fig. 4). For the MLIP trained on the biased dataset, we choose a larger value of $W_{F}/W_{E}$, which is necessary to reach a reasonable accuracy on forces. This comes at the price of sacrificing the accuracy on the energy, a direct consequence of the SCE. The weights used for each MLIP are summarized in Tab. 1. While the values of the best relative training weights $W_{F}/W_{E}$ and $W_{P}/W_{E}$ are model dependent, we highlight that this procedure is generally applicable to any other system.
@@
We run MD simulations at three different densities and temperature conditions close to the position of the deuterium Hugoniot curve given by the VMC-MLIP of Ref. Tenti2024. In particular, we considered thermodynamic conditions that span both the molecular liquid and the atomic one. Because of the $\Delta$-learning approach adopted, the resulting efficiency of these calculations is the same as a DFT Born-Oppeheimer MD simulation. Indeed, the energy, ionic forces, and pressure of the system are calculated at each step with DFT, and the corresponding corrections are then added using the MLIP. Notice how the DFT baseline can be, in principle, replaced by a faster potential (e.g., using an MLIP trained on DFT data Tenti2025), improving the overall efficiency. Further details regarding the computational setup of the MD simulations are given in App. A.4.
A comparison of the equilibrium pressure during the dynamics for three different temperatures is reported in Tab. 2. Notice that, in principle, the reference dataset and the corrected one can give different results. Indeed, the optimization of the antisymmetric part of the WF not only improves the consistency of forces and pressure, but can also modify the PES of the system. Nevertheless, if compared with the biased model, the MLIP trained on the corrected data gives results that are consistently much closer to the reference ones. This suggests that the main discrepancy between the biased and reference models comes from the SCE, and that the SCE correction improves considerably the physical description of the system. We observe a maximum discrepancy of $\sim 2$ GPa at $r_{s}=2.02$ and $T=6000$ K between the biased and corrected results, while the corrected and reference ones are in statistical accordance within a joint error bar of 0.2 GPa at the same thermodynamic conditions. Remarkably, this state is very close to the Hugoniot compressibility peak, i.e., the maximum in the density reached by the shocked state along the Hugoniot curve. This thermodynamic point is very sensitive to changes in the underlying equation of state, and a pressure shift as the one observed here between the corrected and biased models has a significant impact in the final position of the Hugoniot Clay2019. Similar conclusions can be reached by looking at the radial distribution function $g(r)$ (Fig. 5). Analogously to the pressure analysis, the $g(r)$ of the corrected model is compatible with the one given by the reference one. Also for this quantity the maximum discrepancy between the corrected and biased models is observed close to the compressibility peak at $T=6000$ K, which corresponds to the onset of the the molecular-to-atomic transition along the Hugoniot curve. These results further highlight how SCE forces and pressure corrections significantly improve the physical outcome of the resulting MLIPs.
@@ 4 Conclusions
Second, we showed how an easy-to-evaluate bias correction, previously developed in Ref. Nakano2024, can be applied to remove the SCE from the training set, and improve the accuracy of the resulting MLIPs.
@@
Our results reveal the importance of appropriately curing forces and pressures when they are computed with partially optimized WF in QMC and subsequently used to train MLIPs. The MLIPs quality degradation due to the SCE plaguing these VMC datasets is an effect that has been overlooked in previous applications. This could be even more relevant for heavier elements, since the SCE usually grows with the number of valence electrons Tiihonen2021. Moreover, the SCE correction scheme is general and applicable to systems more complex than hydrogen, because it is based on cheaper DFT calculations. We have also shown that some numerical instabilities coming from the finite difference approximation used to evaluate the SCE correction can be cured by an appropriate cutoff in the corresponding error bars. The development of a more efficient scheme for evaluating the SCE correction based on linear response theory is currently in progress. The availability of an SCE correction framework used in combination with other approaches, such as $\Delta$-learning, to further increase the precision of the final models, opens the way to the efficient generation of accurate MLIP based on QMC electronic structure calculations.
@@ A.1 Kernel regression models parameters
In this Appendix we provide further details on the hyperparameters and specific kernels used for the models described in Sec. 3.2. The KRR models use a kernel as the one described in Ref. Tirelli2022. Each environment is described by a density function written as a sum of Gaussian functions (with spread $\epsilon$) multiplied by a localized function $f_{c}$ depending on a cutoff radius $r_{c}$, i.e.,
@@
For the local energy (Eq. (4)) we also added an additional pair-wise term as done in Ref. Tenti2024, using cubic spline functions Xie2023 defined on $10$ equally spaced grid points between $0.3$ Bohr and $5.5$ Bohr.